OpenStudy (anonymous):

use logarithmic differentiation to find the derivative of the function: y=(cos 2x)^x. For an answer i got: (cos2x)^x[(x)(-2tan(x))+ln(cos2x)(1/x)] but it was wrong.

OpenStudy (experimentx):

take log on both sides, 1/y* dy/dx = d(x * log cos(2x))/dx dy/dx = y * d(x * log cos(2x))/dx dy/dx = (cos 2x)^x * d(x * log cos(2x))/dx

OpenStudy (anonymous):

y=(cos 2x)^x. Take ln on both sides lny = ln (cos 2x)^x ln y = x ln cos 2x diff. both side w.r.t. x (1/y)(y') = (1/cos 2x)(-sin 2x) (2) y' = (1/cos 2x)(-sin 2x) (2) (y) =(1/cos 2x)(-sin 2x) (2) (cos 2x)^x = (-2tan2x)(cos 2x)^x

OpenStudy (experimentx):

and that would be (log cos 2x -2 x tan2x )

OpenStudy (anonymous):

Sorry, I have some amendment

OpenStudy (anonymous):

ln y = x ln cos 2x diff. both side w.r.t. x (1/y)(y') = x(1/cos 2x)(-sin 2x) (2) + ln cos 2x y' =x (1/cos 2x)(-sin 2x) (2) (y) + ln cos 2x (y) =(x(1/cos 2x)(-sin 2x) (2) +ln cos 2x )(cos 2x)^x =( x(-2tan2x)(cos 2x)^x +ln cos 2x) (cos 2x)^x

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

answer was (cos 2x)^x (ln cos 2x-2x tan 2x)